OpenStudy (anonymous):

PLEASE HELP!!! How can long division and synthetic division be used to find zeros of polynomial functions, please show one example.

OpenStudy (anonymous):

@Loser66 @SolomonZelman

OpenStudy (loser66):

theoretical question?

OpenStudy (anonymous):

Well its a question on this pre-test and I have to show an example of how

OpenStudy (loser66):

If you have to give out the example by yourself, you should go back ward to make sure that you get the right answer. For example, let do (x+3)(x-2) (x+1) expand them, what do you have? don't forget, this is the BACKWARD process. I want to guide you how to get the correct answer without confusing.

OpenStudy (anonymous):

can you help me work it out? its a pre-test so I have never done it before, sorry

OpenStudy (loser66):

you expand my expression first. give me the expansion expression, then I guide you step by step

OpenStudy (loser66):

(x+3)(x-2)(x+1)= ?

OpenStudy (anonymous):

Would that be = (x+4)(x-2)?

OpenStudy (loser66):

ok, take smaller step (x+3)(x-2) =?

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

?

OpenStudy (loser66):